OpenStudy (anonymous):

Solve. One point for showing the steps to solve each equation and one point for the correct solution. 3m - 5 = 7m - 21

OpenStudy (anonymous):

transpose -21 to LHS as 21 and 3m to RHS as -3m to get -5 + 21 = 7m-3m 16 = 4m dividing both sides by 4 16/4 = 4m/4 4 = m

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

welcome :)